# A DNN Biophysics Model with Topological and Electrostatic Features

### v2

@@ 3 Results
3.1 Specification of the machine learning model
We provide more details of the machine learning model with focus on the neural network and its related specification. The predictive model developed in this study is designed to predict solvation and Coulomb energies based on a combination of topological and electrostatic features. Prior to model training, outliers are identified and removed using the interquartile range (IQR) method. All input features and target labels are standardized to have zero mean and unit variance using the StandardScaler() utility from sklearn.preprocessing. The inverse transformations are applied to the output, returning values with physical meanings e.g. energies in kcal/mol during post-processing for evaluation purposes. The network architecture of the model as shown in Fig. 5 is implemented using the Keras Sequential API with two distinct branches. The first branch as color-coded in blue is a one-dimensional convolutional neural network (CNN) processing the topological features. The CNN branch has three convolution-pooling layers with (128, 64, 64) filters and ($\tanh$, relu, $\tanh$) activation function respectively. The 3 by 3 kernels are used. This CNN branch is responsible for capturing hierarchical and spatially correlated patterns in the topological data. The output of the CNN is flattened and passed through a subsequent Dense layer of 64 units. The other branch is the electrostatic features into two fully connected hidden dense layers with (128, 64) units respectively as shown in orange. After independent processing, the outputs from both branches are concatenated and passed through additional three hidden dense layers with (32, 16, 8) units as shown in purple to produce the final regression output of energy as shown in red.
To enhance generalization and mitigate overfitting, batch normalization is applied after intermediate layers, and dropout is employed to randomly deactivate neurons during training. The model is trained for 500 epochs using the Adam optimizer with a learning rate of 0.0001 and a batch size of 32. The mean squared error (MSE) is used as a loss function. Model architecture and hyperparameters are selected through empirical tuning and informed by previous studies employing Hyperopt [19, 64]. The dataset as described below was split randomly into 80% for training and 20% for testing, and the model was trained with 5-fold cross-validation as well.
3.2 Data
The datasets are sourced from the PDBbind database [65]. In particular, we categorize used data into two datasets as described below. The Amber ff14SB general force field is applied for the atomic van der Waals radii and partial charges when the PDB2PQR [57] software is used.
Dataset 1: These data are from the refined set of the v2018 release [65]. An interquartile range method is used to remove outliers based on $E_{\text{coul}}$, resulting in 4,658 complexes used in our simulation. We also we computed the electrostatic solvation free energies $E_{\text{solv}}$ of the proteins, as part of the 4,658 protein-ligand complexes. All the $E_{\text{solv}}$ values are computed using the 2nd order accurate Matched Interface and Boundary Poisson-Boltzmann (MIBPB) solver [20] at mesh size h=0.2h=0.2Å.
Dataset 2: These data are from the the v2020 release, which has a general set with 14,127 complexes and refined set with 5,316 complexes. We computed the Coulomb energies $E_{\text{coul}}$ of the all proteins using the Treecode [45]. An interquartile range method is used to remove outliers based on $E_{\text{coul}}$, resulting in 17,007 complexes used in our simulation.
3.3 Research Purposes and Metrics
We here provide some simulation results to demonstrate the performance of our model. These results serve three research purposes.
Purpose 1: Training on larger data set receives better performance.
Purpose 2: When electrostatic features are considered, performance of model improves when when treecode level $L$, Taylor expansion order $p$, and their associated number of features $N_{f}$ increases.
Purpose 3: Using both electrostatic features and topological features outperforms using only one kind of these features.
Metrics: We use three common metrics to measure the performance of the model as defined below. In all these metrics, we use $y_{i}$ to denote actual value, $\hat{y}_{i}$ for predicted value, $\bar{y}$ for mean of actual values, $\bar{\hat{y}}$ for mean of predicted values.
MAPE: The mean average percentage errors, which measures prediction accuracy as a percentage. Not suitable if the actual values contain zero or values close to zero.
$\text{MAPE}=\frac{1}{n}\sum_{i=1}^{n}\left|\frac{y_{i}-\hat{y}_{i}}{y_{i}}\right|\times 100\%$
MSE: The mean-squared error, which measures the average absolute difference between predicted and actual values.
$\text{MSE}=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}$
$R^{2}$, which indicates the proportion of variance in the dependent variable that is predictable from the independent variables.
$R^{2}=1-\frac{\text{SS}_{\text{res}}}{\text{SS}_{\text{tot}}}$
3.4 Performance
@@
We next provide simulation results using different dataset and models. On reporting and explaining the performance of each combination of model and dataset, we focus on the achievement of the three aforementioned purposes.
@@
$E_{\text{coul}}$ computed using electrostatic features only on Dataset 1 and Dataset 2
@@
We first investigate the effectiveness of electrostatic features in the prediction of the Coulombic energy $E_{\text{coul}}$. Note $E_{\text{coul}}$ is determined by the location and charges of the particles, which also determines the moments of the multipoles as the electrostatic features.
@@
Table 1: $E_{\text{coul}}$ predicted using Electrostatic Features only on Dataset 1 and Dataset 2; D2-D1 indicates the difference of corresponding values of MSE and $R^{2}$ from Dataset 1 and Dataset 2.
$p$ $L$ $N_{f}(p,L)$ Dataset 1 Dataset 2 D2-D1
@@
Figure 6: (a) $R^{2}$ of model model performance against number of features; (b) scatter plot for Dataset 1 ($p=4$, $L=1$ with total 315 features); (c) scatter plot for Dataset 2 ($p=1$, $L=2$ with total 292 features)
@@
Table 1 reports the $E_{\text{coul}}$ prediction using electrostatic only on Dataset 1 and Dataset 2. The first important pattern we can observe from the table is that the MSE and $R^{2}$ obtained from Dataset 2 are essentially better than those obtained from Dataset 1 as the differences of MSE (D2-D1) are almost all negative and the difference of $R^{2}$ (D2-D1) is almost all positive. This pattern validates Purpose 2 indicating advantage of using larger dataset, particularly at capturing more overall variance in the target. We also noticed the difference of MAPE(D2-D1) is almost all positive, which apparently contradicts the results from MSE and $R^{2}$. In fact, this gives more information about the data distribution, i.e. the data in Dataset 2 but not in Dataset 1 might contain more smaller target values, yielding large percentage errors.
@@
Also from Table 1, we can see on Dataset 1, MSE decreases and $R^{2}$ increase with the increment of $p$ and $L$, reaching optimized value at $p=4$ and $L=1$ with 315 features. Similarly on Dataset 2, MSE decreases and $R^{2}$ increases with the increment of $p$ and $L$ too, reaching optimized value at $p=1$ and $L=2$ with 292 features. This pattern validates the Purpose 1 showing the advantage of using more features. However, we should aware that optimized $p$ and $L$ are affected by the size of the training dataset. When the number of features are larger than number of samples, increasing number of features might bring negative impact such as overfitting.
@@
We plot $R^{2}$ in Fig. 6(a) with horizontal axis as number of features $N_{f}$ from the third column of Table 1. The validation of Purpose 1 is shown since blue squares are lower than the red ones. The validation of Purpose 2 is also shown since markers of the same color contain an increasing pattern as number of features increase until there are too many features. We also plot the best cases on using both Datasets in Fig. 6(b,c) and we can see the distribution of the data points are essentially around the $\hat{y}=y$ line.
$E_{\text{coul}}$ computed using e-features and both features on Dataset 2
@@
Table 2: $E_{\text{coul}}$ computed using e-features and both features on Dataset 2; the MSE and $R^{2}$ for using Topological Features only is 0.039 and 0.962 respectively.
$p$ $L$ $N(p,L)$ e-features both features difference
MSE $R^{2}$ MSE $R^{2}$ MSE $R^{2}$
1 0 4 0.682 0.331 0.039 0.961 -0.643 0.631
2 0 10 0.351 0.655 0.033 0.967 -0.318 0.312
3 0 20 0.252 0.753 0.044 0.956 -0.208 0.203
4 0 35 0.270 0.735 0.030 0.970 -0.240 0.235
0 1 9 0.663 0.349 0.030 0.970 -0.634 0.621
1 1 36 0.235 0.769 0.040 0.960 -0.195 0.191
2 1 90 0.154 0.848 0.044 0.956 -0.110 0.108
3 1 180 0.197 0.806 0.037 0.963 -0.160 0.156
4 1 315 0.224 0.780 0.045 0.955 -0.179 0.175
0 2 73 0.337 0.669 0.048 0.952 -0.289 0.283
1 2 292 0.144 0.859 0.062 0.939 -0.082 0.080
2 2 730 0.170 0.833 0.051 0.949 -0.119 0.116
3 2 1460 0.162 0.841 0.032 0.969 -0.131 0.128
4 2 2555 0.180 0.823 0.039 0.961 -0.141 0.138
0 3 585 0.287 0.719 0.047 0.953 -0.240 0.235
1 3 2340 0.255 0.750 0.050 0.951 -0.205 0.201
2 3 5850 0.215 0.789 0.042 0.958 -0.172 0.169
3 3 11700 0.269 0.736 0.032 0.968 -0.236 0.232
4 3 20475 0.241 0.764 0.037 0.964 -0.204 0.200
We next investigate the effectiveness of combined features. This investigation has two parts. The first part is using $E_{\text{coul}}$ as labels with Dataset 2 and the second part is using $E_{\text{solv}}$ as labels with Dateset 1 considering the availability of labels and the size of the datasets. The MSE and $R^{2}$ for different models are shown in Table 2. Note the MSE and $R^{2}$ for using topological features only is 0.039 and 0.962 respectively, which is already very effective. If we see the difference column, we can see that adding topological features fundamentally improves the model using electrostatic features only by decreasing MSE and increase $R^{2}$. This partially validates purpose 2. Note we did not report MAPE result here as it shows similar pattern as MSE does. If we consider topological features first, adding electrostatic features seems only occasionally improve the results. This is because the results from using topological features are too good for the effectiveness of electrostatic features to kicked in. Our next test case shows when the topological features are relatively less effective, adding the electrostatic features can significantly improve the performance. Before we move to that case, let’s linger on Fig. 7, which is the scatter plot of predicted values vs true values for using topological features, electrostatic feature and combined feature, for the observation that the improvement from (b) to (c) are very obvious while the improvement from (a) to (c), although not that significant, is still discernible.
@@
Figure 7: scatter plots: (a) topological features only, (b) e-features only ($p=1$, $L=2$ with total 292 features), (c) both features ($p=4$, $L=0$ with total 35 features).
@@
$E_{\text{solv}}$ computed using e-feature and both features on Dataset 1
@@
As mentioned earlier, computing electrostatic solvation energies accurately is time consuming. Here we computed the solvation energies for the proteins from Dataset 1 using the MIBPB [20] solver. These solvation energies are used as labels for the neural network training and testing. A similar simulation using Dataset 1 are conducted and the results are reported in Table 3. Form the table, we can see that the improvement on using both features from using topological features only or from using electrostatic features are significant. If we focus on the difference column (both features - e-features only), we can see adding topological features completely improve the model using electrostatic features only by decreasing MSE and increase $R^{2}$. Meanwhile, the MSE and $R^{2}$ in the columns of using both features are obviously better than using topological feature only, which has MSE = 0.150 and $R^{2}$= 0.826. These results validate purpose 3, showing the advantage of using topological and electrostatic features combined. Again, we did not report MAPE result here as it shows similar pattern as MSE does.
@@
Table 3: $E_{\text{solv}}$ computed using e-features and both features on Dataset 1; the MSE and $R^{2}$ for using Topological Features only is 0.150 and 0.826 respectively.
$p$ $L$ $N(p,L)$ e-features both features difference
MSE $R^{2}$ MSE $R^{2}$ MSE $R^{2}$
1 0 4 0.345 0.604 0.063 0.926 -0.282 0.323
2 0 10 0.271 0.690 0.090 0.895 -0.180 0.206
3 0 20 0.135 0.845 0.064 0.926 -0.071 0.081
4 0 35 0.121 0.861 0.065 0.925 -0.056 0.064
0 1 9 0.447 0.487 0.114 0.867 -0.333 0.380
1 1 36 0.148 0.830 0.077 0.911 -0.072 0.081
2 1 90 0.095 0.891 0.084 0.903 -0.012 0.012
3 1 180 0.111 0.873 0.105 0.878 -0.006 0.005
4 1 315 0.128 0.853 0.097 0.887 -0.031 0.034
0 2 73 0.214 0.754 0.092 0.892 -0.122 0.138
1 2 292 0.169 0.806 0.127 0.852 -0.041 0.045
2 2 730 0.164 0.812 0.128 0.851 -0.036 0.039
3 2 1460 0.212 0.757 0.138 0.839 -0.074 0.082
4 2 2555 0.218 0.750 0.165 0.808 -0.053 0.058
0 3 585 0.336 0.615 0.148 0.827 -0.188 0.213
1 3 2340 0.399 0.542 0.218 0.747 -0.181 0.205
2 3 5850 0.390 0.553 0.171 0.801 -0.219 0.249
3 3 11700 0.469 0.461 0.195 0.773 -0.275 0.312
4 3 20475 0.365 0.581 0.242 0.718 -0.123 0.137
The optimized case for using e-feature only happens at $p=1$ and $L=1$ with 90 features. Adding topological features improve the performance. We provide the scatter plot between reference values and predicted values for using topological features only, electrostatic features only, and both features in Fig. 8. The improvements of using two kind of features as shown in (c) can be seen as compared with topological features only in (a) and electrostatic features only in (b).

### v3

@@ 3 Results
We next provide simulation results using different datasets and models. Following that, additional comparisons with baseline models and different grouping strategies are provided to justify the learned model.
3.1 Model Performance
@@
In reporting and explaining the performance of each combination of model and dataset, we focus on achieving the three aforementioned purposes.
[…]
@@
3.1.1 $E_{\text{coul}}$ predicted using electrostatic features only on Dataset 1 and Dataset 2
@@
We first investigate the effectiveness of electrostatic features in predicting the Coulombic energy $E_{\text{coul}}$. Note that $E_{\text{coul}}$ is determined by the locations and charges of the particles, which also determine the moments of the multipoles as electrostatic features.
@@
Table 1: $E_{\text{coul}}$ (kcal/mol) prediction results using electrostatic features only on Dataset 1 and Dataset 2. The D2–D1 columns report the differences (Dataset 2 - Dataset 1) for the corresponding metrics.
$p$ $L$ $N_{f}(p,L)$ Dataset 1 Dataset 2 D2–D1
@@
Figure 8: (a) $R^{2}$ of model performance for models trained on the two datasets against number of features; (b) scatter plot for Dataset 1 ($p=3$, $L=1$ with total 180 features); (c) scatter plot for Dataset 2 ($p=3$, $L=2$ with total 1460 features)
@@
Table 1 reports the $E_{\text{coul}}$ predictions obtained using electrostatic features only on Dataset 1 and Dataset 2. The first important pattern observed from the table is that the MSE and $R^{2}$ obtained from Dataset 2 are generally better than those from Dataset 1, as the differences in MSE (D2-D1) are almost all negative, while the difference in $R^{2}$ (D2-D1) is almost all positive. This pattern validates Purpose 2, indicating the advantage of using a larger dataset, particularly in capturing more overall variance in the target. We also notice that the differences in MAPE(D2-D1) are almost all positive, apparently contradicting the trends observed for MSE and $R^{2}$. In fact, this behavior provides additional insight into the data distribution: Dataset 2, but not Dataset 1, contains more samples with smaller target values as shown in Fig. 7, leading to larger percentage errors.
@@
From Table 1, we also observe that in Dataset 1, the MSE decreases and $R^{2}$ increases as $p$ and $L$ increase, reaching an optimal value at $p=3$ and $L=1$ with 180 features. Similarly, in Dataset 2, the MSE decreases and $R^{2}$ increases with increasing $p$ and $L$, reaching an optimal value at $p=3$ and $L=2$ with 1460 features. This pattern validates Purpose 1, demonstrating the advantage of using more features. However, we should be aware that the optimal choices of $p$ and $L$ are influenced by the size of the training dataset. When the number of features exceeds the number of samples, increasing the number of features may have a negative impact, such as overfitting.
@@
We plot the $R^{2}$ values in Fig. 8(a), where the horizontal axis represents the number of features $N_{f}$ given in the third column of Table 1. The validation of Purpose 1 is evident, as the blue squares are consistently lower than the red ones. The validation of Purpose 2 is also observed, as markers of the same color exhibit an increasing trend with the number of features until the feature set becomes excessively large. We also plot the best-performing cases for both datasets in Fig. 8(b) and (c), where the reference-predicted values are distributed closely around the $\hat{y}=y$ line.
3.1.2 $E_{\text{coul}}$ predicted using e-features and both features on Dataset 2
@@
Table 2: $E_{\text{coul}}$ (kcal/mol) prediction results using electrostatic features and combined features on Dataset 2. The corresponding MSE, MAPE, and $R^{2}$ values obtained using Topological Features only are 0.030, 0.071, and 0.970, respectively.
$p$ $L$ $N_{f}(p,L)$ e-features only both features difference
MSE MAPE $R^{2}$ MSE MAPE $R^{2}$ MSE MAPE $R^{2}$
1 0 4 0.682 0.531 0.323 0.028 0.076 0.972 -0.654 -0.454 0.649
2 0 10 0.340 0.331 0.663 0.035 0.072 0.966 -0.305 -0.258 0.303
3 0 20 0.236 0.284 0.767 0.031 0.064 0.969 -0.205 -0.220 0.202
4 0 35 0.249 0.321 0.753 0.027 0.082 0.973 -0.222 -0.239 0.221
0 1 9 0.633 0.498 0.372 0.029 0.067 0.971 -0.604 -0.431 0.599
1 1 36 0.241 0.273 0.761 0.030 0.068 0.971 -0.211 -0.205 0.210
2 1 90 0.156 0.230 0.845 0.029 0.063 0.971 -0.127 -0.167 0.126
3 1 180 0.163 0.237 0.838 0.029 0.067 0.971 -0.134 -0.170 0.133
4 1 315 0.165 0.234 0.837 0.024 0.073 0.976 -0.141 -0.162 0.140
0 2 73 0.322 0.324 0.681 0.031 0.069 0.969 -0.291 -0.255 0.289
1 2 292 0.127 0.206 0.874 0.035 0.131 0.965 -0.092 -0.075 0.092
2 2 730 0.133 0.208 0.868 0.033 0.109 0.968 -0.101 -0.099 0.100
3 2 1460 0.123 0.195 0.880 0.033 0.092 0.967 -0.089 -0.103 0.087
4 2 2555 0.145 0.229 0.857 0.026 0.067 0.974 -0.119 -0.162 0.117
0 3 585 0.265 0.279 0.737 0.031 0.071 0.969 -0.234 -0.209 0.232
1 3 2340 0.253 0.236 0.752 0.034 0.077 0.967 -0.219 -0.159 0.215
2 3 5850 0.217 0.252 0.784 0.031 0.075 0.969 -0.186 -0.177 0.185
3 3 11700 0.209 0.261 0.792 0.029 0.083 0.971 -0.181 -0.178 0.179
4 3 20475 0.223 0.267 0.779 0.023 0.074 0.978 -0.200 -0.193 0.199
We next investigate the effectiveness of the combined feature representation. This analysis consists of two parts. In the first part, $E_{\text{coul}}$ is used as the label on Dataset 2, while in the second part, $E_{\text{solv}}$ is used as the label on Dataset 1, taking into account the size of the datasets and the availability of labels.
@@
Figure 9: Scatter plots for the prediction of $E_{\text{coul}}$ on Dataset 2: (a) topological features only, (b) e-features only ($p=3$, $L=2$ with total 1460 features), (c) both features ($p=4$, $L=1$ with total 315 features).
The MSE, MAPE, and $R^{2}$ values for the different models are reported in Table 2. Using topological features alone yields MSE, MAPE, and $R^{2}$ values of 0.030, 0.071, and 0.970, respectively, indicating strong predictive accuracy. Moreover, the difference column clearly shows that incorporating topological features into the electrostatics-only model consistently reduces MSE and MAPE while increasing $R^{2}$. These results partially support Purpose 2.
If we consider topological features first, it appears that adding electrostatic features does not significantly improve the results. This is because the performance obtained using topological features is already very strong, leaving limited room for the effectiveness of electrostatic features to manifest. Our next test case shows that when the topological features are relatively less effective, adding electrostatic features can significantly improve performance. Before moving to that case, we briefly examine Fig. 9, which shows scatter plots of predicted values versus true values using topological features, electrostatic features, and combined features. The improvement from (b) to (c) is very evident, while no obvious improvement is observed from (a) to (c).
@@
3.1.3 $E_{\text{solv}}$ predicted using e-features only and both features on Dataset 1
@@
As mentioned earlier, computing electrostatic solvation energies accurately is time consuming. Here, we use the MIBPB solver Chen:2011 to compute the solvation energies of proteins in Dataset 1, which are then used as labels for model training and testing. A similar simulation using Dataset 1 is conducted, and the results are reported in Table 3. From the table, we can see that the improvement obtained by using both features, compared to using only topological features or only electrostatic features, is significant. If we focus on the difference column, we see that adding topological features clearly improves the electrostatics-only model by reducing the MSE and MAPE while increasing the $R^{2}$ score. In addition, the results obtained using both features are noticeably better than those obtained using topological features alone, which yield MSE = 0.150, MAPE = 0.153, and $R^{2}$ = 0.826. These results validate purpose 3, and demonstrate the advantage of combining topological and electrostatic features.
@@
Table 3: $E_{\text{solv}}$(kcal/mol) computed using e-features and both features on Dataset 1; the MSE, MAPE and $R^{2}$ for using Topological Features only is 0.150, 0.153 and 0.826 respectively.
$p$ $L$ $N_{f}(p,L)$ e-features only both features difference
MSE MAPE $R^{2}$ MSE MAPE $R^{2}$ MSE MAPE $R^{2}$
1 0 4 0.345 0.243 0.604 0.063 0.108 0.926 -0.282 -0.136 0.323
2 0 10 0.271 0.256 0.690 0.090 0.094 0.895 -0.180 -0.162 0.206
3 0 20 0.135 0.142 0.845 0.064 0.081 0.926 -0.071 -0.061 0.081
4 0 35 0.121 0.138 0.861 0.065 0.083 0.925 -0.056 -0.055 0.064
0 1 9 0.447 0.318 0.487 0.114 0.169 0.867 -0.333 -0.149 0.380
1 1 36 0.148 0.161 0.830 0.077 0.104 0.911 -0.072 -0.057 0.081
2 1 90 0.095 0.130 0.891 0.084 0.098 0.903 -0.012 -0.032 0.012
3 1 180 0.111 0.138 0.873 0.105 0.103 0.878 -0.006 -0.036 0.005
4 1 315 0.128 0.144 0.853 0.097 0.105 0.887 -0.031 -0.039 0.034
0 2 73 0.214 0.205 0.754 0.092 0.116 0.892 -0.122 -0.089 0.138
1 2 292 0.169 0.156 0.806 0.127 0.123 0.852 -0.041 -0.033 0.045
2 2 730 0.164 0.156 0.812 0.128 0.116 0.851 -0.036 -0.039 0.039
3 2 1460 0.212 0.168 0.757 0.138 0.124 0.839 -0.074 -0.045 0.082
4 2 2555 0.218 0.168 0.750 0.165 0.128 0.808 -0.053 -0.040 0.058
0 3 585 0.336 0.211 0.615 0.148 0.135 0.827 -0.188 -0.076 0.213
1 3 2340 0.399 0.216 0.542 0.218 0.143 0.747 -0.181 -0.073 0.205
2 3 5850 0.390 0.202 0.553 0.171 0.139 0.801 -0.219 -0.063 0.249
3 3 11700 0.469 0.242 0.461 0.195 0.145 0.773 -0.275 -0.097 0.312
4 3 20475 0.365 0.210 0.581 0.242 0.156 0.718 -0.123 -0.053 0.137
The optimized case for the electrostatic-only model occurs at $p=2$ and $L=1$ with 90 features. Adding topological features further improves the performance, achieving an MSE of 0.064, a MAPE of 0.081, and an $R^{2}$ value of 0.926 using only 20 features at $p=3$ and $L=0$. Figure 10 presents scatter plots for the optimized cases, comparing reference and predicted values for models using topological features only, electrostatic features only, and the combined feature set. The benefits of combining both types of features, as illustrated in (c), can be clearly observed when compared with using only topological features in (a) and only electrostatic features in (b).